Shift Summary 10 Aug., 1999 evening.

Nickolas Solomey (solomey@hep.uchicago.edu)
Tue, 10 Aug 1999 23:47:35 -0500

Shift Summary for Evening shift, 10 Aug 99

On shift: Herman and Nick

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: "shutdown in progress" all day

=======================================================================
Tapes taken to FCC: QKa24x (for a=C,S,K and x=6,7,8) and QKB086
========================================================================
Runs ended this shift: None to speak of that people need to know about.
========================================================================
Accesses during shift: All day
========================================================================
Overview of shift:
Scheduled accelerator shutdown in progress. KTeV hive is
buzzing with activity. Several accesses to conduct repairs finished,
and hall secured.
========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
15:30 Herman from new shift arrives, finds a lot of activity.
16:00 Nick arrives half hour late due to Cascade muon decay PRL
meeting with Godfathers, Herman says the activity is
dwindling.
16:20 HV for CsI channel changed from 1112 to 1327 V for crystal
number 1691 by Jason Hamm.
16:40 Visual inspection of hall and helium bags completed
Crash buttons on the analysis magnet are left engaged
until Search & Secure request:(D. Northacker, Herman, Patricia)
17:30 D. Northacker is repairing the air duct to the Etot power supply
rack on the west side column plenum. This repair amounts to
replacing inefficient air duct material, requiring under floor
access.
17:45 Nick takes MINOS visitors for a tour.
18:20 Dave Northacker completes work.
18:32 Etot power restored (Jason Hamm)
Crash buttons on analysis magnet pulled out (HBW)
18:37 Control Room was called to request search and secure. They are
short of personnel, but will honor the request in a couple of
hours.
21:14 Search and secure team plus Nick start enclosure interlock
procedure.
21:32 Nick survives his first search and secure.
23:15 Bob T and Rick K make an access to work on a DYC in ADC crate 8
that might be causing a problem hanging the DAQ.
23:50 New shift arrives and we negotiate leaving a few min early.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey